OpenStudy (anonymous):

I need help converting 4.8 lbs to grams. I have to use decimal analysis

OpenStudy (aaronq):

what is the conversion factor between pounds and grams? it's called \(\sf dimensional\) analysis, btw.

OpenStudy (anonymous):

There are 453.5 grams in a pound. And sorry about that, I'm running on cold medicine, not a good combination with chemistry.

OpenStudy (aaronq):

sorry to hear that, being sick is never a good combination with studying lol so 1 lb = 453.5 g make this into a fraction, so that you cancel out the units you dont want. \(\sf 4.8 lbs *\dfrac{453.5~g}{1~lb}=4.8*453.5~\dfrac{\cancel{lbs}*g}{\cancel{lbs}}=4.8*453.5~g\)

OpenStudy (anonymous):

Thanks. So it's 2176.5g

OpenStudy (aaronq):

if the conversion factor you provided is correct, then heck yes.

OpenStudy (anonymous):

I googled it to mkae sure I had the correct factor so yay :)

OpenStudy (aaronq):

nice! though i hope you understood the process

OpenStudy (aaronq):

of cancelling out units

OpenStudy (anonymous):

Yeah, I got that part. I just wasn't completely sure if the problem was set up the same as all the other ones I'd done before because I was converting to a smaller unit